OpenStudy (anonymous):

Prove (sinsquaredx+4sinx+3)/cossquaredx=(sinx+3)/1-sinx

OpenStudy (blurbendy):

is the left side all over cos^2x ?

OpenStudy (blurbendy):

or just the 3

OpenStudy (anonymous):

all of it

OpenStudy (anonymous):

there i edited it

sam (.sam.):

\[\frac{\sin^2x+4sinx+3}{\cos^2x}=\frac{sinx+3}{1-sinx}?\]

sam (.sam.):

What's the question?

OpenStudy (anonymous):

yeah thats how it is, i dont know how to start

sam (.sam.):

Prove? Simplify? Solve?

OpenStudy (anonymous):

prove

sam (.sam.):

First you gotta factor the numerator

sam (.sam.):

what do you get?

OpenStudy (anonymous):

um (sinx+1)(sinx+3)?

sam (.sam.):

Good, you noticed that the RHS got all the terms in sine, so we're gonna change the denominator to sine, by using \(cos^2x=1-sin^2x\) \[\frac{(sinx+1)(sinx+3)}{1-\sin^2x}\] The denominator can be factored as well, try factoring it

OpenStudy (anonymous):

how would i factor that?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

obviously

OpenStudy (anonymous):

(1+sin)(1-sin)

sam (.sam.):

Yes

sam (.sam.):

Then just cancel and you got it

OpenStudy (anonymous):

aha thanks !

sam (.sam.):

yw :)